OpenStudy (wade123):

Eliminate the parameter. x = t2 + 2, y = t2 - 4

OpenStudy (aum):

Just subtract the two equations and t^2 will get eliminated.

OpenStudy (jdoe0001):

\(\bf x=t^2+2\implies x-2=t^2\implies \sqrt{x-2}={\color{blue}{ t}} \\ \quad \\ y={\color{blue}{ t}}^2-4\implies y=?\)

OpenStudy (wade123):

y = x - 6, x ≥ 1? @jdoe0001

OpenStudy (jdoe0001):

notice above... use substitution

OpenStudy (wade123):

oh the x is squared?

OpenStudy (wade123):

y = x2 - 6

OpenStudy (jdoe0001):

hmmm ohh shoot.. ahemm you're correct dohh .. yes

OpenStudy (wade123):

wait i was right before? it would be y = x - 6, x ≥ 1?

OpenStudy (jdoe0001):

\(\bf x=t^2+2\implies x-2=t^2\implies \sqrt{x-2}={\color{blue}{ t}} \\ \quad \\ y={\color{blue}{ t}}^2-4\implies y=(\sqrt{x-2})^2-4\to x-2-4\to x-6\) so long \(\bf x > 1\)

OpenStudy (wade123):

so yay i was right(: thanks

OpenStudy (jdoe0001):

the domain constraint will apply to the rooted version thus if x is equals to 1, that gives \(\bf \sqrt{1-2}\to \sqrt{-1}\) but if x is greater than 1, the radicand will be 0 or a positive value

OpenStudy (wade123):

so i was right..

OpenStudy (jdoe0001):

yes, just keep in mind the domain is x > 1